# Wasserstein Transfer Learning

### v1

@@ 6 Real-world Applications
During the 2005–2006 NHANES cycle, participants aged 6 and older wore an ActiGraph 7164 accelerometer on their right hip for seven days, recording physical activity intensity in 1-minute epochs. Participants were instructed to remove the device during water-based activities and sleep. The device measured counts per minute (CPM), ranging from 00 to $32767$, capturing variations in activity levels throughout the monitoring period.
@@
To evaluate the WaTL, we set White, Mexican Americans, and other Hispanic individuals as sources and Black as target. For females, the source data include $1308$ White people, $884$ Mexican Americans, and $108$ Other Hispanic individuals. For males, the source data include $1232$ White people, $805$ Mexican Americans, and $92$ Other Hispanic individuals. We set $200$ Black participants as the target data for both genders. During evaluating, we perform five-fold cross-validation, using four folds for training and one for testing, cycling through each fold. For comparison, we also apply local Fréchet regression using only the target data. The results, summarized in Figure 2 (a), show that WaTL improves performance over local Fréchet regression for both females and males, demonstrating its ability to leverage information from other participants to enhance modeling for black participants.
@@
[…]
@@ 7 Conclusion
We introduce Wasserstein transfer learning and its adaptive variant for scenarios where the informative set is unknown, addressing the challenges posed by the lack of linear operations in the Wasserstein space. By leveraging the Wasserstein metric, the proposed algorithm accounts for the non-Euclidean structure of distributional outputs, ensuring compatibility with the geometry of the Wasserstein space. Supported by rigorous theoretical guarantees, the framework demonstrates improved estimation performance compared to methods that rely solely on target data.
This paper focuses on univariate distributions, which arise frequently in real-world applications such as the NHANES and human mortality studies discussed in Section 6. The proposed framework, however, is not limited to the univariate case and can be extended to multivariate distributions by incorporating metrics such as the Sinkhorn (Cuturi 2013) or sliced Wasserstein distance (Bonneel et al. 2015). Extending the methodology to higher dimensions presents an exciting future direction and would involve addressing both computational and theoretical challenges specific to high-dimensional Wasserstein spaces.
Our theoretical analysis assumes that each probability distribution is fully observed, consistent with much of the existing literature on distributional data analysis (Petersen & Müller 2019, Chen et al. 2023). However, this assumption could be unrealistic in practical applications, where only independent samples from the underlying distributions are available. This limitation can be addressed by replacing the unobservable distributions with empirical measures constructed from these samples (Zhou & Müller 2024). The derived convergence rates can be extended to incorporate an additional term reflecting the number of independent samples per distribution, providing a straightforward way to account for this practical consideration.
[…]
We further assess WaTL using the age-at-death distributions from 162 countries in 2015, compiled from the United Nations Databases [^2] https://data.un.org/ and the UN World Population Prospects 2022333https://population.un.org/wpp/Download. The dataset provides country-level age-specific death counts, which we convert into smooth age-at-death densities using local linear smoothing; see Figure 3 (a) for an illustration. For this analysis, we define the 24 developed countries as the target site and the remaining 138 developing countries as the source site.
Given the limited sample size, we input the entire dataset into the model without performing cross-validation. Country-level age-at-death distributions are shaped by a range of factors, including economic conditions, healthcare systems, and social and environmental contexts. Among these, we select population density and gross domestic product (GDP) per capita as key predictive covariates. The results, shown in Figure 3 (b), demonstrate that the proposed WaTL method successfully transfers knowledge across these demographically diverse regions, effectively addressing the socioeconomic disparities between the source and target domains.
Figure 3: (a) Age-at-death densities of developed and developing countries; (b) Root mean squared prediction risk (RMSPR) of WaTL and Only Target methods for human mortality data.

### v2

@@ 6 Real-world Applications
During the 2005–2006 NHANES cycle, participants aged 6 and older wore an ActiGraph 7164 accelerometer on their right hip for seven days, recording physical activity intensity in 1-minute epochs. Participants were instructed to remove the device during water-based activities and sleep. The device measured counts per minute (CPM), ranging from 0 to $32767$, capturing variations in activity levels throughout the monitoring period.
@@
To evaluate the WaTL, we set White, Mexican Americans, and other Hispanic individuals as sources and Black as the target. For females, the source data include $1308$ White people, $884$ Mexican Americans, and $108$ Other Hispanic individuals. For males, the source data include $1232$ White participants, $805$ Mexican Americans, and $92$ Other Hispanic individuals. We set $200$ Black participants as the target data for both genders. During evaluating, we perform five-fold cross-validation, using four folds for training and one for testing, cycling through each fold. For comparison, we also apply local Fréchet regression using only the target data. The results, summarized in Figure 2(a), show that WaTL improves performance over local Fréchet regression for both females and males, demonstrating its ability to leverage information from other demographic groups to enhance modeling for Black participants.
@@
To further illustrate the effectiveness of WaTL, we visualize the cumulative distribution function of physical activity levels for one female participant and one male participant in Figure 2(b), along with estimates from WaTL and local Fréchet regression, using only the target data. The results indicate that WaTL provides a better fit to the true distribution, outperforming the estimate obtained using only the target data. We further evaluate the robustness of WaTL on a human mortality dataset in Appendix 8, where WaTL continues to demonstrate strong performance.
@@ 7 Conclusion
We introduce Wasserstein transfer learning and its adaptive variant for scenarios where the informative set is unknown, addressing the challenges posed by the lack of linear operations in the Wasserstein space. By leveraging the Wasserstein metric, the proposed algorithm accounts for the non-Euclidean structure of distributional outputs, ensuring compatibility with the intrinsic geometry of the Wasserstein space. Supported by rigorous theoretical guarantees, the framework demonstrates improved estimation performance compared to methods that rely solely on target data.
This paper focuses on univariate distributions, which arise frequently in real-world applications such as the NHANES and human mortality studies discussed in Section 6. The proposed framework, however, is not limited to the univariate case and can be extended to multivariate distributions by incorporating metrics such as the Sinkhorn (Cuturi 2013) or sliced Wasserstein distance (Bonneel et al. 2015). Extending the methodology to higher dimensions is an exciting future direction that entails addressing both computational and theoretical challenges specific to high-dimensional Wasserstein spaces.
Acknowledgments
We thank the Area Chair and the reviewers for their constructive feedback. Doudou Zhou was supported by the NUS Start-Up Grant (A-0009985-00-00) and the MOE AcRF Tier 1 Grant (A-8003569-00-00).